OpenStudy (artemisxrpg):

Find one angle theta that satisfies each of the following:

OpenStudy (artemisxrpg):

cos\[\cos \theta=\sin(\theta/4 + 3 degree)\]

OpenStudy (anonymous):

just convert both sides in terms of cos angle or sin angle and then equate the angles

OpenStudy (anonymous):

like r.gupta said \[\cos \theta = \sin (\theta + \frac{\pi}{2})\] \[-\sin \theta = \cos (\theta + \frac{\pi}{2})\] ^^

OpenStudy (anonymous):

those identities are in radians, however. Keep that in mind.